OpenStudy (anonymous):

Could someone please help me with these? I hate converting! Jane poured 12 cups of water into a rectangular container measuring 12 inches by 8 inches by 9 centimeters. [1 cup = 14.44 cubic inches; 1 inch = 2.54 cm] Part A: What is the maximum volume of water that the rectangular container can hold? Show your work. (3 points) Part B: How many cubic inches of water was poured into the container? (3 points) Part C: What was the height of the water in the container? Show your work. (4 points) Measurement is given as length by width by height.

hero (hero):

V = lwh

OpenStudy (anonymous):

I'm aware of finding volume, it's the conversion with cubic inches and inches

hero (hero):

But first before solving anything we have to convert centimeters to inches

OpenStudy (anonymous):

9 centimeters = 3.54 inches

hero (hero):

Very good

hero (hero):

So now you can compute the volume of the rectangular container

hero (hero):

Since 1 = 12 w = 8 h = 3.54

hero (hero):

So V = (12)(8)(3.54)

OpenStudy (anonymous):

=339.84

hero (hero):

Don't forget units

hero (hero):

When we multiplied each unit, inches got multiplied three times so 339.84 cubic inches.

hero (hero):

Now it asks how much volume of water was poured into the tank. It gives 1 cup = 14.44 cubic inches All you have to do now is multiply both sides by 12: (12)(1 cup) = 12(14.44) cubic inches

OpenStudy (anonymous):

so 173.28 cubic inches

hero (hero):

Correct

hero (hero):

For part C, in order to find the height of the water, you have to take 173.28 and divide it by the given length and width.

hero (hero):

So \[\frac{\text{Volume of water}}{\text{length} \times \text{width}}\]

OpenStudy (anonymous):

so = 1.805

hero (hero):

Congratulations, you are correct.

hero (hero):

But don't forget units

hero (hero):

The height of the water is 1.805 inches

OpenStudy (anonymous):

You're a lifesaver, thank you so much!

OpenStudy (anonymous):

@vinurules wheres the answer?